2.25 moles of an ideal gas at 35.6oC expands isothermally from an initial volume of 26.0 dm3

to a

final volume of 70.0 dm3

. Calculate w for this process (a) for expansion against a constant external

pressure of 1.00 x 105 Pa and (b) for a reversible expansion.


1
Expert's answer
2021-09-04T02:28:47-0400

70 - 26 = 44 dm3

= 2.25 × 308.6 × 44 × 1 × 105

= 3.0551 × 109 j